A flour mill is considering buying a new jumbo sifter, which would have an installed cost of $80,000. The new one would replace the existing sifter that was purchased 3 years ago at an installed cost of $60,000. If the company moves forward with the replacement, it could sell the old sifter for $19,000. Purchasing the new sifter would result in the company's current assets increasing by $10,000 and current liabilities increasing by $8,000. The company uses the 5-year MACRS table for depreciation and is taxed at 21%. a) What is the accumulated depreciation of the old sifter? b) What is the current book value of the old sifter? c) What is the amount of depreciation recapture/recovery? d) What is the tax on the sale of the old sifter? e) What are the after-tax proceeds from the sale of the old sifter? f) What is the change in Net Working Capital? g) What is the initial investment for the project?
Answer #1

Accumulated depreciation using MACRS 5 years class = Cost*(20%+32%+19.20%)

= $60,000*71.2%

= $42,720

b)Current book value = Cost - Accumulated depreciation

= 60,000-42,720

= $17,280

c)gain on sale or depreciation recovery = Selling price - book value

= 19000-17,280

= $1,720

d)Tax = 1720*21% = $361.2

e)After tax proceeds = 19000-361.2

= $18,638.8

f)Change in NWC = Increase in Current Assets - Increase in Current Liabilities

= 10,000-8000

= $2000

g)Initial Investment = Cost of new + Increase in NWC - after tax proceeds of old

= 80,000+2000-18638.8

= $63,361.2